OpenStudy (anonymous):

Choose the equation that represents the graph below. y = x − 3 y = −x + 3 y = −x − 3 y = x + 3

OpenStudy (anonymous):

here is the graph

OpenStudy (anonymous):

4th choice

OpenStudy (anonymous):

The equation is given by x/x-intercept + y/y-intercept = 1. So, x/-3 + y/3 = 1 , y-x = 3 , y = x+3

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

Yw.